Ticket Number: HEA-NOC/20070412-13 Ticket Status: UPDATE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20070412 16:16 UTC+1 Problem Start: 20070412 16:11 UTC+1
Ticket Update: 20070614 14:50 UTC+1 Problem End:
Site/line: Blanch Services Network, CWT Services Network, DNS, Debian
Packages, Listserv, MX backup, NIS, Netflow, Network
Management, Orion Backup server, SmokePing, Web Hosting,
Web-Server, eduroam, hot-standby, mail, mirrors, General
Debian have recently released version 4.0 of their operating system. An
upgrade procedure has been created, and rolling upgrades of Debian servers
will take place over a period of a few weeks.
A variety of servers and services
20070412 16:13 BS opening ticket to track work
20070416 14:28 BS 5 servers updated so far. Interesting issues discovered
on canyonero, the server used to sync mirror content.
20070508 14:19 BS upgrades to 14 hosts have been completed, 25 to go.
byron upgrade scheduled for this evening, meaning that all email services
on byron (including heanet.ie and schools.edu.ie) will be at risk.
20070612 15:14 BS twilight upgraded to sarge - short outage of
OpenGroupware was experienced due to NPTL incompatabilities.
20070612 15:17 BS further work on twilight.heanet.ie scheduled for later
20070614 11:47 BS 20 servers left to do
20070614 12:00 BS upgrades of mokusei, kidman, guzzler, gobbler and kestra
20070614 14:15 BS upgrades completed
20070614 14:34 BS devourer, dosei, hitchcock, kubrick & ironhide next
This ticket can be monitored at http://www.hea.net/tickets/20070412-13
HEAnet Limited. Registered in Ireland, No. 275301.